OpenStudy (anonymous):

2 tanks contain equal amounts of water. They are connected by a pipe and 3000 litres of water is pumped from one tank to the other. One tank then contains 6 times as much water as the other. How many litres of water was there originally in each tank?

OpenStudy (mayankdevnani):

Let's suppose X is the mount of water of one tank,,, X + 3000 = 6(X - 3000)--- equation

OpenStudy (anonymous):

x+3000=6*(x-3000) => x=4200 lt of water was there in each tank initially Verify: 4200-3000=1200 4200+3000-7200

OpenStudy (mayankdevnani):

then substitute value...can you @Genuine

OpenStudy (anonymous):

Oh thank you both for helping me :D

OpenStudy (mayankdevnani):

welcome